OpenStudy (anonymous):

Can someone tell me how to multiply these together. For some reason i forgot 6(x+a)(x-a) + 6(x-a)(x-1) The book says 12x^2-12x I have no idea how they got that. Please explain how the answer was obtained. Do you distribute the 6 to both binomials? and then combine? But that makes no sense

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

first factor out the binomials then multiply the 6 out

OpenStudy (anonymous):

so u foil the binomials into one trem then distribute the 6

OpenStudy (anonymous):

yeah basically

OpenStudy (anonymous):

that doesnt work

OpenStudy (anonymous):

are you sure that's the correct answer? you might want to look back at the answer there should be an a in it or else it doesnt work..

OpenStudy (anonymous):

thats what the book says there is a picture

OpenStudy (anonymous):

(x-a)(x+a) =x^2 - a^2 (x-a)(x-1) = x^2 +a -x -ax

OpenStudy (anonymous):

would u care to explain? that obviously is not a foil

OpenStudy (anonymous):

I know what the answer is but how did they get it

OpenStudy (anonymous):

I know what the answer is but how did they get it

OpenStudy (anonymous):

ah I think either the question or the book's answer is wrong

OpenStudy (anonymous):

i literally cut and pasted it out ot of the book

OpenStudy (anonymous):

lol sorry but that is what I get

OpenStudy (anonymous):

The book is wrong or your problem is copied wrong. Trust me.

OpenStudy (anonymous):

^^^^^^^^^^^

OpenStudy (anonymous):

I second that

OpenStudy (anonymous):

you guys so the picture i i put up right? cuz that is a snip in from the ebook so if thats wrong then the book is wrong because i cut and psted so its not error on my part correct?

OpenStudy (anonymous):

I have a few wrong answers in my math book. it is possible

OpenStudy (anonymous):

thanks all